# LubDubDecoder: Bringing Micro-Mechanical Cardiac Monitoring to Hearables

### v1

@@ 3.3.2. Signal processing
For each region, the average power was computed as $P=\frac{1}{N}\sum\limits_{i=1}^{N}x[i]^{2}$, where $x[i]$ are the amplitudes and $N$ is the number of samples in the region. We denote $P_{signal}$ and $P_{noise}$ as the power within the signal and noise region respectively, $SNR$ is then computed as $SNR=10\log_{10}(\frac{P_{signal}}{P_{noise}})$.
@@ 3.3.3. Model architecture
The encoder employs a two-branch convolutional neural network. The local branch uses dilated 1D convolutions with small kernel sizes (3 $\times$ 1) to capture local temporal patterns within a single cardiac cycle, while the global branch uses large receptive fields (48 $\times$ 1) to extract inter-beat context and periodic structures. The outputs from both branches are concatenated and passed through a temporal self-attention module, which dynamically weighs feature contribution over time. MaxPooling and Dropout layers follow each block for dimensionality reduction and regularization. Residual connections and batch normalization are integrated throughout the architecture to improve training stability and gradient flow.
Following the encoder, additional convolutional and attention layers refine the latent representation. The decoder then mirrors the encoder through a series of upsampling and convolutional layers. Attention modules are integrated throughout the decoder to maintain temporal coherence and selectivity. A final 1D convolution layer that projects the intermediate features into a single channel representing the reconstructed SCG/GCG waveform.
@@ 4.1. Hardware setup
Figure 11. Demographic summary of participants in the human subjects study. Complete demographic data was collected for 12 of 18 participants.
[…]
This figure consists of four bar charts arranged horizontally, presenting the demographic data of the study’s participants. The first bar chart, titled ”Age (years),” shows the distribution of participant ages, with the largest group of 9 participants aged 19.0 to 25.4 years old. The remaining three age groups, 25.4 to 31.8, 31.8 to 38.1, and 63.6 to 70.0, each contain 1 participant. The second chart, ”Gender,” shows that there are 10 male and 2 female participants. The third chart, ”Height (cm),” displays height distribution. There is 1 participants between 160.0 cm and 167.5 cm, 5 participants between 167.5 cm and 175.0 cm, 2 participants between 175.0 cm and 182.5 cm, and 4 participants with a height between 182.5 cm and 190 cm. The fourth chart, ”Weight (kg),” shows the distribution of participant weights. There is 4 participant weighing between 49.9 to 62.9, 5 participants weighing between 62.9 kg and 76.0 kg, 1 participants between 76.0 kg and 89.0 kg, and 2 participants with a weight between 89.0 kg and 102.0 kg.
@@ 4.2. Dataset collection protocol
This study was approved by our Institutional Review Board  (STUDY2025_00000155). All studies complied with relevant ethical regulations. Participants were recruited by word of mouth through our university campus. Written consent was obtained for human subjects participating in the study. Randomization was not applicable and investigators were not blinded.
Participants above the age of 18 were eligible for the study. Exclusionary criteria include pregnant women, prisoners, adults with cognitive impairments, individuals with allergies or sensitivities to adhesives or sensors, individuals with implanted medical devices, such as pacemakers, individuals with significant mobility impairments. Individuals were screened via inclusion and exclusion criteria prior to the study. We recruited 18 adults with a female-to-male ratio of 0.39, of whom 12 had complete demographic records: mean age 27 ± 14 years, height 177 ± 9 cm, weight 71 ± 15 kg. All of whom were healthy and without history of cardiovascular disease (Fig. 11).
We collected data in a quiet room from using the hearables to record ear-based cardiac sounds, and the smartphone IMU for SCG and GCG data. To synchronize the cardiac signals between the hearable and the phone, we tapped both together twice. Participants then remained in a static seated position for 10 minutes as data was recorded. After each measurement session, we aligned the recordings based on the tap events. We collected a total of 570 minutes (9.5 hours) of synchronized data from the participants. In a deployment of our system, this synchronization could similarly be achieved by having the smartphone transmit an inaudible acoustic chirp (18–20 kHz) that is simultaneously received by the smartphone and hearable microphones.
For all 18 participants, we collected recordings from the in-ear speaker of wired earbuds, and for 15 of them, also from the in-ear microphone of wireless earbuds. In addition, for a subset of 3 participants recordings were obtained across all five hearables (Fig. 5).
@@ 5.2. Waveform reconstruction performance
Figure 12. Waveform reconstruction performance across 18 subjects. Results are shown for within-user and cross-user evaluations.
[…]
This figure is composed of two bar charts, one above the other, that display the waveform reconstruction performance for two different signals across 18 subjects. The top chart, labeled (a) SCG similarity, shows the Pearson Correlation for the Seismocardiogram (SCG) signal. The bottom chart, labeled (b) GCG similarity, shows the Pearson Correlation for the Gyrocardiogram (GCG) signal. In both charts, each of the 18 subjects is represented by a pair of bars. The first bar in each pair, a stippled orange, represents the Within-User reconstruction performance, while the second bar, a solid navy blue, represents the Cross-User performance. Error bars are included to show standard deviation. Horizontal dashed lines on each chart indicate the Mean Within-User and Mean Cross-User correlation values.
@@
Within-user. We first evaluate the reconstruction performance of SCG and GCG waveforms under a within-user setting, where both training and testing are performed on data from the same participant. When using the in-ear speaker of wired earbuds, across 18 subjects, the reconstructed signals closely match the ground truth, achieving average Pearson correlation scores of 0.95 $\pm$ 0.04 for SCG and 0.94 $\pm$ 0.04 for GCG (Fig. 12a).
@@
Cross-user. We then perform cross-user evaluation by training a model on all but one user, and fine-tuning it on five calibration cardiac cycles from the held-out user. Under this protocol, the system using the in-ear speaker achieves mean Pearson correlation scores of 0.88 $\pm$ 0.07 for SCG and 0.89 $\pm$ 0.05 for GCG (Fig. 12b). Similarly, when using the in-ear microphone, the system achieved scores of 0.84 $\pm$ 0.13 for SCG and 0.85 $\pm$ 0.09 for GCG.
@@ 5.3. Timing accuracy of micro-cardiac events
[…]
This figure consists of four Cumulative Distribution Function (CDF) plots arranged in a two-by-two grid, illustrating the timing accuracy of five micro-cardiac events for both SCG and GCG signals. The top row, labeled ”(a) Within-user,” shows results for the same user, while the bottom row, labeled ”(b) Cross-user,” shows results across different users. The left column of plots is for SCG signals, and the right is for GCG signals. All plots share the same axes: ”Timing Error (ms)” on the x-axis and ”CDF” on the y-axis. Each plot contains five lines, one for each event, with a legend listing the median and 95th percentile timing errors. For the Within-user SCG plot, the median and 95th percentile errors are: MC (2ms, 20ms), IM (2ms, 10ms), AO (0ms, 4ms), MA (2ms, 20ms), and RE (2ms, 18ms). For the Within-user GCG plot, they are: MC (2ms, 26ms), IM (2ms, 8ms), AO (0ms, 4ms), MA (2ms, 18ms), and RE (4ms, 18ms). For the Cross-user SCG plot, they are: MC (4ms, 20ms), IM (2ms, 18ms), AO (0ms, 4ms), MA (4ms, 22ms), and RE (4ms, 26ms). Finally, for the Cross-user GCG plot, they are: MC (6ms, 32ms), IM (2ms, 16ms), AO (2ms, 4ms), MA (4ms, 16ms), and RE (4ms, 30ms).
@@
We next evaluate the timing error of key micro-cardiac fiducial points (MC, IM, AO, MA, RE) between the reconstructed and reference measured waveform.
@@
Within-user. Fig. 15a shows that, for SCG signals, the median absolute errors (95th percentile) were 2 ms (20 ms) for MC, 2 ms (10 ms) for IM, 0 ms (2 ms) for AO, 2 ms (20 ms) for MA, and 2 ms (16 ms) for RE. For GCG signals, the corresponding errors were 2 ms (26 ms) for MC, 2 ms (10 ms) for IM, 0 ms (4 ms) for AO, 2 ms (18 ms) for MA, and 4 ms (18 ms) for RE. These values translate to relative median timing errors of 0.0–0.5% and 95th percentile errors of 0.5 to 3.3%, assuming an 800 ms cardiac cycle typical of a healthy adult heartbeat at 75 bpm.
@@
Cross-user. As illustrated in Fig. 15b, the median absolute errors (95th percentile) for SCG were 4 ms (20 ms) for MC, 2 ms (18 ms) for IM, 0 ms (4 ms) for AO, 4 ms (22 ms) for MA, and 4 ms (26 ms) for RE. For GCG, the corresponding values were 6 ms (32 ms) for MC, 2 ms (16 ms) for IM, 2 ms (4 ms) for AO, 4 ms (16 ms) for MA, and 4 ms (30 ms) for RE. These correspond to relative median timing errors of 0.3–0.8% and 95th percentile errors of 0.5–4.0%.
@@
We note that prior work (Ha et al., 2020) using mmWave for SCG reconstruction obtains similar median timing errors of 0.3–1.6%. Furthermore, the variability of fiducial point timing as mentioned in our study is often higher within a single session ranging from 1—8 ms for SCG and GCG (Table 3).
@@
Cross-device. We next evaluated the timing accuracy of reconstructed SCG fiducial points after applying our cross-device normalization. As shown in Fig. 14b, the system achieved median absolute errors (95th percentile errors) of 2 ms (20 ms) for MC, 2 ms (20 ms) for IM, 0 ms (6 ms) for AO, 4 ms (18 ms) for MA, and 4 ms (28 ms) for RE. These correspond to relative median errors in the range of 0.0–0.5% and 95th percentile errors of 0.75–3.5% of the cardiac cycle, which are comparable to our within-user and cross-user performance results.
@@ 5.4. Subgroup analysis
Figure 16. Subgroup analysis on cross-user waveform reconstruction performance.
[…]
This figure contains two adjacent bar charts, each presenting a subgroup analysis of the cross-user waveform reconstruction performance. The y-axis for both charts is labeled Pearson Correlation, ranging from 0.0 to 1.0. The left chart displays performance based on body mass index (BMI) subgroups, while the right chart shows performance based on gender. In the left chart, the performance for the underweight subgroup (n=1) shows a mean Pearson correlation of 0.95 for SCG and 0.91 for GCG. The normal weight subgroup (n=9) shows a mean correlation of 0.86 with a standard deviation of 0.20 for SCG, and 0.89 with a standard deviation of 0.13 for GCG. The overweight subgroup (n=1) shows a mean correlation of 0.85 for SCG and 0.86 for GCG. Finally, the obese subgroup (n=1) has a mean correlation of 0.90 for SCG and 0.89 for GCG. In the right chart, the male subgroup (n=13) has a mean correlation of 0.91 with a standard deviation of 0.10 for SCG, and 0.92 with a standard deviation of 0.08 for GCG. The female subgroup (n=5) shows a mean correlation of 0.87 with a standard deviation of 0.18 for SCG, and 0.88 with a standard deviation of 0.13 for GCG. Error bars on all bars represent the standard deviation, and horizontal dashed lines indicate the overall mean Pearson correlation for SCG and GCG across all groups. For both figures, SCG refers to the color of dark blue, while GCG refers to the color of orange.
@@
Body mass index. In our study, 8% ($n=1$) of subjects were underweight ($BMI<18.5$), 75% ($n=9$) had normal weight ($18.5\leq BMI<24.9$), 8% ($n=1$) were overweight ($25.0\leq BMI<29.9$), and 8% ($n=1$) were obese ($BMI\geq 30$). For the underweight group, the system achieved average reconstruction correlation of $0.95\pm 0.03$ (SCG) and $0.91\pm 0.07$ (GCG). In the normal weight group, similarities were $0.86\pm 0.20$ (SCG) and $0.89\pm 0.13$ (GCG). Among overweight subjects, results were $0.85\pm 0.13$ (SCG) and $0.86\pm 0.10$ (GCG). Finally, in the obese group, the system reached $0.90\pm 0.09$ (SCG) and $0.89\pm 0.10$ (GCG).
Sex. In our study, 28% ($n=5$) of subjects were female and 72% ($n=13$) of subjects were male. When comparing system performance, female subjects had an SCG reconstruction similarity of $0.91\pm 0.10$ and a GCG similarity of $0.92\pm 0.08$, while male subjects had an SCG similarity of $0.87\pm 0.18$ and a GCG similarity of $0.88\pm 0.13$.

### v2

@@ 3.4.2. Signal processing
Multi-cycle training. While single-cycle inputs allow the model to learn detailed intra-beat morphology, they provide limited information about the inter-beat features across consecutive heartbeats. To address this, we incorporate a multi-cycle training strategy in which the continuous signal is segmented into 3-second windows that contain multiple cardiac cycles. These longer segments enable the model to capture inter-beat dynamics especially when the user’s heart rate changes across time, such as after exercise, or during sleeping. Combined with single-cycle training, this hybrid approach allows the model to learn both fine-grained intra-beat features and broader inter-beat patterns, improving overall reconstruction fidelity.
To help the model learn inter-beat patterns across a wide range of heart rates, we apply a stretching-based data augmentation strategy (Ha et al., 2020). Specifically, each segment is time-stretched by factors ranging from 0.8 to 1.3, effectively simulating slower and faster heart rates. This augmentation improves the model’s ability to generalize to different heart rates during real-world use.
@@
For each region, the average power was computed as
$P=\frac{1}{N}\sum_{i=1}^{N}x[i]^{2},$
where $x[i]$ are the amplitudes and $N$ is the number of samples in the region. We denote $P_{\mathrm{signal}}$ and $P_{\mathrm{noise}}$ as the power within the signal and noise regions, respectively. The SNR is then computed as
$\mathrm{SNR}=10\log_{10}\!\left(\frac{P_{\mathrm{signal}}}{P_{\mathrm{noise}}}\right).$
@@ 3.4.3. Model architecture
The encoder employs a two-branch convolutional neural network with progressively increasing channel dimensionality ($64\to 128\to 256\to 512$). The local branch uses dilated 1D convolutions with small kernel sizes (3 $\times$ 1) to capture fine-grained temporal patterns within individual cardiac cycles, while the global branch employs larger receptive fields (48 $\times$ 1) to model inter-beat context and periodic structure. The outputs from both branches are concatenated and refined through a temporal self-attention module that dynamically reweights feature contributions over time. Each dual-path block is followed by MaxPooling for temporal downsampling and Dropout for regularization. Residual connections and batch normalization layers are integrated throughout to stabilize training and ensure smooth gradient flow.
Following the encoder, three additional convolutional layers with 512 channels and batch normalization further refine the latent representation, preserving temporal continuity while enhancing local feature selectivity. The decoder mirrors the encoder through four upsampling blocks, each consisting of an upsampling layer (scale factor of 2) followed by a convolutional layer (kernel size 3), batch normalization, and LeakyReLU activation. Channel dimensionality is progressively reduced from $512\to 256\to 12\to 64\to 32$ to reconstruct the temporal resolution of the original signal. Attention modules are integrated within both the encoder’s dual-path blocks and decoder to enhance temporal coherence and focus on salient cardiac events. Finally, a 1D convolution layer projects the intermediate feature maps into a single output channel representing the reconstructed SCG or GCG waveform.
@@ 4.1. Hardware setup
Figure 10. Demographic summary of participants in the human subjects study. Complete demographic data was collected for 19 of 25 participants. Gender data was collected for all 25 participants
Four histogram panels showing demographic data for our study. Panel 1 (Age) shows the distribution of 19 participants across age ranges from 20 to 70+ years, with the highest count of approximately 7 participants in the 19-25 age range, and smaller counts distributed across older age groups. Panel 2 (Gender) displays a bar chart of 25 participants with 18 males and 7 females. Panel 3 (Height) shows the distribution of 19 participants’ heights in centimeters, ranging from 160 to 190 cm. Panel 4 (Weight) shows the distribution of 19 participants’ weights in kilograms, ranging from 50 to 100 kg.
@@ 4.2. Dataset collection protocol
This study was approved by our Institutional Review Board
(STUDY2025_00000155). All studies complied with relevant ethical regulations. Participants were recruited by word of mouth through our university campus. Informed written consent was obtained for human subjects participating in the study before any study procedures began. Randomization was not applicable and investigators were not blinded.
Participants above the age of 18 were eligible for the study. Exclusionary criteria include pregnant women, prisoners, adults with cognitive impairments, individuals with allergies or sensitivities to adhesives or sensors, individuals with implanted medical devices, such as pacemakers, individuals with significant mobility impairments. Individuals were screened via inclusion and exclusion criteria prior to the study. We recruited 25 adults with a female-to-male ratio of 0.39, of whom 19 had complete demographic records: mean age 36.4 ± 17.9 years, height 176.6 ± 7.8 cm, weight 72.0 ± 13.7 kg. All of whom were healthy and without history of cardiovascular disease (Fig. 10).
We collected data in a quiet room from using the hearables to record ear-based cardiac sounds, and the smartphone IMU for SCG and GCG data. To synchronize the cardiac signals between the hearable and the phone, we tapped both together twice. Participants then remained in a static seated position for 10 minutes as data was recorded. After each measurement session, we aligned the recordings based on the tap events. We collected a total of 710 minutes (11.8 hours) of synchronized data from the participants. In a deployment of our system, this synchronization could similarly be achieved by having the smartphone transmit an inaudible acoustic chirp (18–20 kHz) that is simultaneously received by the smartphone and hearable microphones.
For all 25 participants, we collected recordings from the in-ear speaker of wired earbuds, and for 15 of them, also from the in-ear microphone of wireless earbuds. In addition, for a subset of 3 participants recordings were obtained across all five hearables (Fig. 4).
@@ 5.2. Waveform reconstruction performance
Figure 11. Waveform reconstruction performance across 25 subjects. Results are shown for within-user and cross-user evaluations.
This figure is composed of two bar charts, one above the other, that display the waveform reconstruction performance for two different signals across 25 subjects. The top chart, labeled (a) SCG similarity, shows the Pearson Correlation for the Seismocardiogram (SCG) signal. The bottom chart, labeled (b) GCG similarity, shows the Pearson Correlation for the Gyrocardiogram (GCG) signal. In both charts, each of the 25 subjects is represented by a pair of bars. The first bar in each pair, a stippled orange, represents the Within-User reconstruction performance, while the second bar, a solid navy blue, represents the Cross-User performance. Error bars are included to show standard deviation. Horizontal dashed lines on each chart indicate the Mean Within-User and Mean Cross-User correlation values.
@@
Within-user. We first evaluate the reconstruction performance of SCG and GCG waveforms under a within-user setting, where both training and testing are performed on data from the same participant. When using the in-ear speaker of wired earbuds, across 25 subjects, the reconstructed signals closely match the ground truth, achieving average Pearson correlation scores of 0.94 $\pm$ 0.04 for SCG and 0.93 $\pm$ 0.04 for GCG (Fig. 11a).
@@
Cross-user. We then perform cross-user evaluation by training a model on all but one user, and fine-tuning it on five calibration cardiac cycles from the held-out user. Under this protocol, the system using the in-ear speaker achieves mean Pearson correlation scores of 0.88 $\pm$ 0.07 for SCG and 0.88 $\pm$ 0.05 for GCG (Fig. 11b). Similarly, when using the in-ear microphone, the system achieved scores of 0.84 $\pm$ 0.13 for SCG and 0.85 $\pm$ 0.09 for GCG.
@@ 5.3. Timing accuracy of micro-cardiac events
Four cumulative distribution function (CDF) plots showing timing accuracy of micro-cardiac events for SCG (seismocardiography) and GCG (gyrocardiography) measurements collected on in-ear speaker earbuds. The figure is organized in two rows: (a) Within-user performance and (b) Cross-user performance. Each row contains left panels showing SCG timing error and right panels showing GCG timing error. In panel (a) within-user SCG (left), five overlapping curves represent different cardiac events: MC (mitral valve closure), IM (isovolumic moment), AO (aortic valve opening), MA (mitral valve anterior), and RE (rapid ejection). All curves rise steeply near 0ms timing error, with AO showing the best performance (median=0ms, 95th=2ms) and RE showing slightly lower performance (median=2ms, 95th=20ms). The within-user GCG panel (right) shows similar patterns with median timing errors between 0-4ms and 95th percentiles between 2-18ms. In panel (b) cross-user evaluation, both SCG (left) and GCG (right) show slightly degraded but still strong performance. SCG median timing errors range from 0-4ms with 95th percentiles between 2-26ms, while GCG shows medians of 0-4ms with 95th percentiles between 4-30ms. All curves demonstrate that the system maintains sub-30ms timing accuracy at the 95th percentile across different cardiac event types in both within-user and cross-user scenarios, with AO and IM events showing the most precise timing detection.
Within-user (SCG)
MC IM AO MA RE
Median 2.0[2.0,2.0] 2.0[2.0,2.0] 0.0[0.0,0.0] 2.0[2.0,2.0] 2.0[2.0,4.0]
95th 18.0[16.0,20.0] 8.0[6.0,12.0] 2.0[2.0,2.0] 18.0[16.0,22.0] 20.0[15.1,27.1]
Within-user (GCG)
Median 4.0[2.0,4.0] 2.0[2.0,2.0] 0.0[0.0,0.0] 2.0[2.0,2.0] 4.0[2.0,4.0]
95th 30.6[26.0,37.2] 14.0[10.0,20.0] 2.0[2.0,4.0] 16.0[14.0,20.0] 18.0[16.0,20.6]
Cross-user (SCG)
MC IM AO MA RE
Median 4.0[4.0,4.0] 2.0[2.0,2.0] 0.0[0.0,0.0] 4.0[4.0,4.0] 4.0[4.0,4.0]
95th 26.0[24.0,26.0] 24.0[22.0,24.0] 2.0[2.0,2.0] 22.0[20.0,22.0] 22.0[22.0,24.0]
Cross-user (GCG)
Median 4.0[4.0,4.0] 2.0[2.0,4.0] 2.0[2.0,2.0] 4.0[4.0,4.0] 4.0[4.0,4.0]
95th 28.0[26.0,30.0] 18.0[18.0,20.0] 4.0[4.0,6.0] 16.0[16.0,18.0] 30.0[29.8,34.0]
Cross-device (SCG)
MC IM AO MA RE
Median 2.0[2.0,4.0] 2.0[2.0,2.0] 0.0[0.0,0.0] 4.0[2.0,4.0] 4.0[4.0,4.0]
95th 20.0[18.0,22.0] 20.0[16.0,22.5] 6.0[4.5,8.0] 18.0[16.0,18.0] 28.0[26.0,30.0]
@@
Table 4. Timing error (ms) of micro-cardiac fiducial points between reconstructed and reference waveforms. Values are median and 95th percentile absolute error with 95% confidence intervals.
We next evaluate the timing error of key micro-cardiac fiducial points (MC, IM, AO, MA, RE) between the reconstructed and reference waveforms. For each point, Table 4 reports the median and 95th percentile absolute errors with 95% confidence intervals (CI), along with a significance test across participants.
@@
Within-user. Fig. 14a summarizes within-user performance, and Table 4 reports the corresponding timing errors for SCG and GCG. Using a Wilcoxon signed-rank test across participants, we found no significant difference in the timing errors for any fiducial point (SCG: all $p>0.14$; GCG: all $p>0.38$). These values translate to relative median timing errors of 0.0–0.5% and 95th percentile errors of 0.5–3.9%, assuming an 800 ms cardiac cycle typical of a healthy adult heartbeat at 75 bpm.
@@
Cross-user. As illustrated in Fig. 14b, Table 4 reports cross-user timing errors for SCG and GCG. We again found no significant differences across participants (SCG: all $p>0.23$; GCG: all $p>0.31$). These correspond to relative median timing errors of 0.0–0.5% and 95th percentile errors of 0.5–3.8%.
@@
We note that prior work (Ha et al., 2020) using mmWave for SCG reconstruction obtains similar median timing errors of 0.3–1.6%. Furthermore, the within-session variability of fiducial point timing is often higher, ranging from 1–8 ms for SCG and GCG (Table 3).
@@
Cross-device. We evaluated the timing accuracy of reconstructed SCG fiducial points after applying cross-device normalization. As shown in Fig. 13b and Table 4, the system achieved comparable timing accuracy to within-user and cross-user settings. Using a Wilcoxon signed-rank test across participants, we found no significant difference in timing errors for any fiducial point (all $p>0.16$). These correspond to relative median errors of 0.0–0.5% and 95th percentile errors of 0.75–3.5% of the cardiac cycle.
@@ 5.4. Subgroup analysis
Figure 15. Subgroup analysis on cross-user waveform reconstruction performance by (a) BMI, (b) gender, (c) age.
This figure contains two adjacent bar charts, each presenting a subgroup analysis of the cross-user waveform reconstruction performance. The y-axis for both charts is labeled Pearson Correlation, ranging from 0.0 to 1.0. The left chart displays performance based on body mass index (BMI) subgroups, while the right chart shows performance based on gender. In the left chart, the performance for the underweight subgroup (n=2) shows a mean Pearson correlation of 0.93 for SCG and 0.82 for GCG. The normal weight subgroup (n=12) shows a mean correlation of 0.85 with a standard deviation of 0.08 for SCG, and 0.88 with a standard deviation of 0.06 for GCG. The overweight subgroup (n=4) shows a mean correlation of 0.87 for SCG and 0.87 for GCG. Finally, the obese subgroup (n=1) has a mean correlation of 0.87 for SCG and 0.88 for GCG. In the right chart, the male subgroup (n=18) has a mean correlation of 0.86 with a standard deviation of 0.07 for SCG, and 0.88 with a standard deviation of 0.04 for GCG. The female subgroup (n=7) shows a mean correlation of 0.90 with a standard deviation of 0.03 for SCG, and 0.88 with a standard deviation of 0.07 for GCG. Error bars on all bars represent the standard deviation, and horizontal dashed lines indicate the overall mean Pearson correlation for SCG and GCG across all groups. For both figures, SCG refers to the color of dark blue, while GCG refers to the color of orange. The figure shows that among the 19 participants whom we have full demographic information in our study, 37% (n=7) of them were among the age of 18-24, 21% (n=4) were between 25-32, 32% (n=6) were 41-55. and 11% (n=2) were 70+. For the 18-24 age group, our system achieved average reconstruction correlation of 0.88 +/- 0.05 (SCG) and 0.89 +/- 0.04 (GCG). In the 25-32 age group, our system achieved reconstruction correlation of 0.87 +/- 0.09 (SCG) and 0.92 +/- 0.03 (GCG). In the 41-55 age group, our system achieved reconstruction correlation of 0.86 +/- 0.04 (SCG) and 0.85 +/- 0.04 (GCG). For the 70+ age group, our system achieved reconstruction correlation of 0.82 +/- 0.09 (SCG) and 0.78 +/- 0.05 (GCG).
@@
Body mass index. Among the 25 participants that participated in our study, only 19 of them shared their BMI data. 11% ($n=2$) of subjects were underweight ($BMI<18.5$), 63% ($n=12$) had normal weight ($18.5\leq BMI<24.9$), 21% ($n=4$) were overweight ($25.0\leq BMI<29.9$), and 5% ($n=1$) were obese ($BMI\geq 30$). For the underweight group, the system achieved average reconstruction correlation of $0.93\pm 0.02$ (SCG) and $0.82\pm 0.09$ (GCG). In the normal weight group, similarities were $0.85\pm 0.08$ (SCG) and $0.88\pm 0.25$ (GCG). Among overweight subjects, results were $0.87\pm 0.09$ (SCG) and $0.87\pm 0.01$ (GCG). Finally, in the obese group, the system reached $0.90\pm 0.09$ (SCG) and $0.89\pm 0.10$ (GCG).
Sex. In our study, 28% ($n=7$) of subjects were female and 72% ($n=18$) of subjects were male. When comparing system performance, female subjects had an SCG reconstruction similarity of $0.90\pm 0.03$ and a GCG similarity of $0.88\pm 0.07$, while male subjects had an SCG similarity of $0.87\pm 0.07$ and a GCG similarity of $0.88\pm 0.05$.
Age. Among the 19 participants whom we have full demographic information in our study, 37% ($n=7$) of them were among the age of 18-24, 21% ($n=4$) were between 25-32, 32% ($n=6$) were 41-55. and 11% ($n=2$) were 70+. For the 18-24 age group, our system achieved average reconstruction correlation of $0.88\pm 0.05$ (SCG) and $0.89\pm 0.04$ (GCG). In the 25-32 age group, our system achieved reconstruction correlation of $0.87\pm 0.09$ (SCG) and $0.92\pm 0.03$ (GCG). In the 41-55 age group, our system achieved reconstruction correlation of $0.86\pm 0.04$ (SCG) and $0.85\pm 0.04$ (GCG). For the 70+ age group, our system achieved reconstruction correlation of $0.82\pm 0.09$ (SCG) and $0.78\pm 0.05$ (GCG). The results show that our system continues to achieve reliable performance in older adults.